hi, i need help on my wrksht. it says to graph these linear equations using 3 points. the equation is 3x-4y=12. please help me and explain how u got this answer.thanks in advance.can u please answer my question before sunday.
Answer by music974(43)

To graph the linear equation you must plug in any three numbers to x or y.The first number can be x=0. 3(0)-4y=12, 0-4y=12, -4y=12.After you do these three steps you then divide each side by -4.Then you would get y=-3.So the first point would be (0,-3).The second number would be y=0. 3x-4(0)=12, 3x-0=12, 3x=12. Now you divide each side by 3 which gives you x=4. Your second point would be (4,0). The third number would be x=8. 3(8)-4y=12, 24-4y=12, -4y=-12. Now you divide each side by -4 which gives you y=3. The third point would be (8,3). The three points would be (0,-3),(4,0),(8,3).